OpenStudy (anonymous):

ok obiviously im trying to figure out all types of questions that im a little iffy on before I take my trig finale. so I have a word problem I don't seem to understand: two radar stations are 50 miles apart. Both are tracking the same plane. at the moment, the angle between station 1 and the plane is 73 degrees, and the angle between station 2 and the plane is 46 degrees how far is the plane from station 2.

OpenStudy (jkristia):

I think you can use law of sin for this. where \[\frac{ a }{ \sin A } = \frac{ b }{ \sin B } = \frac{ c }{ \sin C }\] 

OpenStudy (jkristia):

Oops, I swapped c and C, C is the angle, c is the length. in your case you have A and B, so you can find C.

OpenStudy (anonymous):

so \[A \div \sin(73)= B \div \sin(46)=C \div \sin(x) \] where do I go from there

OpenStudy (jkristia):

So x is 180 - 73 - 46 = 61 so you have the ratio is 50/sin(61) = 57.16. Then you should be able to find the remaining 2 sides

OpenStudy (anonymous):

ok so if I am using angle B in my case it is equal to 46 degrees. and trying to find length a what side would each other length be so I can determine what I use. because I think hypotenuse would be b and adjacent would be C so that would be 50 and then solve for the opposite, but im confused on how I would figure out where to put each side length. in other words. if I wanted to use sin that equals O/H, would that work to solve for the side im looking for or would that solve for the other side then I would need to use coswhich is A/h because then I would only have one more side I was looking for?

OpenStudy (jkristia):

sorry, got interrupted \[\frac{ 50 }{ \sin(61) } = 57.16\] and you want to find length a, so you need sin(A) \[\frac{ a }{ \sin(A) } = 57.16 => a = 57.16 * \sin(A)\] Remember to set you calculator to degree